OpenStudy (cheeks):

If a central angle of measure 30° is subtended by a circular arc of length 6 meters, as is illustrated below, how many meters in length is the radius of the circle?

OpenStudy (campbell_st):

the formula for arc length is simply \[l = \frac{\theta}{360} \times 2\pi r\] you know l = 6 and theta = 30 substituting anf you get \[6 = \frac{30}{360} \times 2\times \pi \times r\] just solve for r